# Improving global awareness of linkset predictions using Cross-Attentive Modulation tokens

### v2

@@ Abstract
Whether it is to predict social network users affinity, molecule chemical bounds, or communication network links, link prediction has been an important field of research for several years. While many machine learning solutions have been proposed to predict single links or sets of independent links, the resolution of a range of problems requires predicting multiple interdependent links. Examples of this type of problems can typically be algorithmic graph problems, which involve finding a set of links satisfying an objective or optimizing some metrics. Using Deep Learning methods as a way to tackle such problems and replace combinatorics or handcrafted heuristics by deep feature learning has been receiving great attention recently, and appears to be a major asset to enable much faster and constant time processing of such tasks, whether they aim at computing strong heuristics using objective-driven embedding processes, or directly provide valid solutions using end to end deep learning methods. Other domains, such as molecule discovery, rely on deep learning techniques to infer novel, plausible graphs that should resemble a collection of dataset graphs and strictly adhere to the same constraints. Most of these techniques rely on the attention mechanism or on Graph Neural Networks (GNNs), which consist in leveraging node-level information exchanges in order to form proper link predictions. Such node-level interactions do not process nodes as an ordered sequence, which would imply some kind of natural ordering of the nodes: they are said to be permutation invariant mechanisms. They are well suited for graph problems, but struggle at providing a global orchestration of the predicted links, which can result in a loss of performance. Some typical issues can be the difficulty to ensure high-level properties such as global connectedness, fixed diameter or to avoid information bottleneck effects such as oversmoothing and oversquashing, which respectively consist in abundant smoothing in dense areas leading to a loss of information and a tendency to exclude isolated nodes from the message passing scheme, and often result in irrelevant, unbalanced link predictions.
To tackle this problem, we hereby present Cross-Attentive Modulation (CAM) tokens, which introduce cross-attentive units used to condition node and edge-level modulations in order to enable context-aware computations that improve the global consistency of the prediction links. We will implement it on a few permutation invariant architectures, and showcase benchmarks that prove the merits of our work.
@@ 1 Introduction
Learning representations on unstructured or semi-structured data has been one of the major deep learning topics throughout the last decade. From images to graphs by way of text data, multiple architectures have been proposed, and most of them rely on token-level (pixel patches, nodes, edges, words…) feature aggregators that are well fitted to the properties and symmetries of the data modality. Training such aggregators allows to capture robust patterns that grant good generalization properties. For instance, ConvNets [1], Transformers [2] or GNNs [3], which respectively deal with images, text and graphs, are probably some the most popular and best examples of architectures designed to deal with unstructured or semi-structured data. Most of these are local or relative, which means that they learn a systematic operator that is applied on each input element independently to allow it to gather some information about the others elements. This way, each input unit progressively gains information about the other units. To do so, ConvNets mainly learn convolution kernels, while Transformers learn a retrieval-like Query-Key mechanism and GNNs typically learn a message passing scheme. This way of processing inputs differs from other traditional Feedforward Neural Networks such as Multi-layer Perceptrons [4], which directly combine the inputs according to learned inter-neuron connection patterns. While this straightforward approach can be perfect for many tasks, it is not well suited for most structured data tasks, where the order of the inputs does not indicate any information about the input. When all inputs follow the same set of computations, regardless their position in the input sequence, they are said to go through a permutation-invariant architecture. One major drawback of such architectures is the lack of global pattern aggregation capabilities of the mechanisms that are used. This is especially the fact when dealing with graphs, where nodes struggle to keep track of long range or graph-level dependencies, and tend to e oversmoothing and oversquashing. The domains of generative graph modeling and multi-link prediction can also particularly suffer from this phenomena, as there is no general mechanism that dictates the global distribution of the generated nodes or links, which makes the inference of valid graphs quite challenging. Our goal is to find, for any input set of nodes $V$, a plausible corresponding set of edges $E$. This is often treated as a probabilistic problem that implies estimating the conditional joint distribution of the edges $E$ knowing $V$, which is a classical formulation for conditional generative models. Several variational methods such as [5] or [6] have been proposed. The first one relies on a node-level sampling procedure that assumes nodes’ independence, the second one uses a graph level latent space that is then decoded into a graph. The main difficulty of such method is that it requires complex matching algorithms to compare the reconstructed nodes and edges with the labeled ones. When the decoding is node-conditioned, the conditioning signal also serves as a node identifier (a reconstructed node to which is concatenated an attribute $a_{k}$ will be compared to the labeled node with attribute $a_{k}$ if the attributes are unique), which can allow avoiding the matching process. Other methods use generative adversarial formulations to generate realistic graphs. In [7], the authors use a unconditional Generative Adversarial Network (GAN) helped by objective-driven Reinforcement Learning, an iterative GAN is proposed in [8] while a temporal sequence GAN is showcased in [9]. Flow-based models, such as autoregressive [10] are also a possible solution for graph generation. More recently, Langevin-esque variational methods have become really popular to generate graphs, in particular Denoising Diffusion Probabilistic Models (DDPM) such as [11] or [12] that respectively provide rotation invariance and discrete, sparsity-encouraging noise schedule. Both use Graph Transformers [13] to denoise at both node and edge levels, as edge attributes are modulated by the nodes’ attention scores, which they also modulate, in an interdependent way. Graph Transformers are quite popular and useful to operate at both node and edge levels, and the attention mechanism is particularly well fitted for graph problems, and grants permutation invariance. As mentioned beforehand, as expressive as it is, the attention mechanism can still struggle to learn graph-level features and is not totally immune to some artifacts, whether by saturating some central nodes with information or creating unwanted high predicted-link-density clusters in areas with an important concentration of nodes. One popular way of improving and providing a form of conditioning to the feature aggregator is to use Feature-wise Linear Modulation (FiLM) [14] [15], which maps a conditional signal to an affine modulation of a feature. For instance, in [12], the authors use FiLM layers to condition the layers on the time step of the diffusion process, to feature some additional graph-level attributes, and encapsulate some global, statistical information about the nodes and edges. In [16], the authors propose using registers, inspired by [17], which consists in adding dummy nodes that can be attended to by the true nodes as supplementary information storage. Using attention on the true tokens to gather general information can also so be seen in other domains such as natural language processing with BERT’s CLS tokens [18], in CrossVIT [19] or in [20] for session-based recommendation. While such method is particularly light, it still has limited expressiveness for graphs and does not provide a flexible mechanism to tune the prediction of the links.
Our work will mainly consist in leveraging flexible and expressive information gathering with cross-attention-based register tokens, and in using this information as a conditioning signal for node-level and edge-level modulations.
We will start by introducing in detail Cross-Attentive Modulation (CAM) tokens, and present some simple models and architectures in which we incorporate the token. A brief introduction to our toy datasets will then be followed by benchmarks comparing the previously mentioned models’ performance with and without CAM tokens.
@@ 2 Our problem
Throughout this work, we will tackle a simple toy problem that typically requires both local and global arbitration: given a set of nodes $V$ described by random 2D coordinates, we wish to predict a minimal set of edges $E$, subject to several constraints. One node can not have more than $k$ edges, edges must respect a range constraint, which implies that their length has to be inferior to $d$. The predicted edges have to minimize the number of connected components, hence, if it is possible regarding the range constraint, the graph has to be connected, if not, the number of connected components has to be as small as possible. This can be viewed as a constrained Minimum Spanning Tree problem with all edges’ weights being equal to 1. It then implies respecting some local constraints and requires some higher level perception of connectivity of the graph, which is typically what we try to improve, and what node-level message passing schemes can struggle to tackle. The aim of such approach is not necessarily to get neural networks to perfectly mimic exact optimization results, but to be able to infer plausible graphs that resemble a dataset of graphs that follow the aforementioned constraints. In order to be valid, the inferred edges $E$ have to: - be individually valid, present similar properties as the dataset edges and, most importantly, respect the range constraints
@@ 5 State of the art generative method: Probabilistic Denoising
Recent works have proven Denoising Diffusion Probabilistic Models (DDPMs) to be particularly effective to tackle the task of generating plausible graphs. In order to approximate the distribution that we assume our data to follow, DDPMs for graphs propose, in the continuous setting, to search for a $p_{\theta}^{t}(\mathbf{G^{t-1}}|\mathbf{G^{t}})$ that approximates a markovian reverse noise process which, contrarily to directly maximizing the likelihood of $p_{\theta}(\mathbf{E}|\mathbf{V})$, is directly tractable. In our case, we want to compute the edges $E$ given the nodes $V$, hence we apply the denoising process on edges only. We would then wish to find $p_{\theta}^{t}(\mathbf{E^{t-1}}|\mathbf{E^{t}},V^{t})=p_{\theta}^{t}(\mathbf{E^{t-1}}|\mathbf{G^{t}})$.
In the discrete setting, which we use as it proved to perform better, we directly model $p_{\theta}^{t}(\mathbf{E^{t=0}}|\mathbf{G^{t}})$. The noise is applied to state transition matrices between the possible discrete categories. Here we consider a binary 1-dimensional problem, this is why our "transition matrix" is then only a scalar $m$ that is correlated with the true distribution of the dataset links.
The noise at step $t$ is then given by $\bar{Q}^{t}=\bar{\alpha}^{t}I+\bar{\beta}^{t}\mathbf{m}$, where $\bar{\alpha}^{t}=\cos({0.5\pi(t/T+s)}/(1+s))^{2}$ with a small $s$ and $\bar{\beta}^{t}=1-\bar{\alpha}^{t}$.
The learning procedure then solely resides in minimizing $l_{CE}(p_{\theta}^{t}(G^{t}),E^{t=0})$.
Even though DDPMs noise and denoise each edge independently, the denoising decisions should depend on the other edges in order to grasp the interdependence between them while iterating towards realistic solutions. Once again, Graph Transformers are a valid architecture to enable node-level edge-conditioned message passing. Each edge can then infer its own value, depending on its initial value and on the information it has gathered through message-passing. Sampling with our DDPM to generate new graphs then iteratively follows
$E^{t-1}\ \mathtt{\sim}\ \prod_{ij}p_{\theta_{ij}}^{t}(e_{ij}^{t-1}|E^{t})q(e_{ij}^{t-1}|e_{ij}=e,e_{ij}^{t}),$ (17)
with $q$ representing the prior distribution of the edges used in the diffusion process (which simply follows the scalar value $m$ in our case) and $p_{\theta}$ being parametrized by a neural network.

### v4

@@ Abstract
This work introduces Cross-Attentive Modulation (CAM) tokens, which are tokens whose initial value is learned, gather information through cross-attention, and modulate the nodes and edges accordingly. These tokens are meant to improve the global awareness of link predictions models which, based on graph neural networks, can struggle to capture graph-level features. This lack of ability to feature high level representations is particularly limiting when predicting multiple or entire sets of links. We implement CAM tokens in a simple attention-based link prediction model and in a graph transformer, which we also use in a denoising diffusion framework. A brief introduction to our toy datasets will then be followed by benchmarks which prove that CAM token improve the performance of the model they supplement and outperform a baseline with diverse statistical graph attributes.
@@ 1 Introduction
Learning representations on unstructured or semi-structured data has been one of the major deep learning topics throughout the last decade. From images to graphs by way of text data, multiple architectures have been proposed, and most of them rely on token-level (pixel patches, nodes, edges, words…) feature aggregators that are well fitted to the properties and symmetries of the data modality. Training such aggregators allows to capture robust patterns that grant good generalization properties. For instance, ConvNets [1], Transformers [2] or GNNs [3], which respectively deal with images, text and graphs, are probably some the most popular and best examples of architectures designed to deal with unstructured or semi-structured data. Most of these are local or relative, which means that they learn a systematic operator that is applied on each input element independently to allow it to gather some information about the others elements. This way, each input unit progressively gains information about the other units. To do so, ConvNets mainly learn convolution kernels, while Transformers learn a retrieval-like Query-Key mechanism and GNNs typically learn a message passing scheme. This way of processing inputs differs from other traditional Feedforward Neural Networks such as Multi-layer Perceptrons [4], which directly combine the inputs according to learned inter-neuron connection patterns. While this straightforward approach can be perfect for many tasks, it is not well suited for most structured data tasks, where the order of the inputs does not indicate any information about the input. When all inputs follow the same set of computations, regardless their position in the input sequence, they are said to go through a permutation-invariant architecture. One major drawback of such architectures is the lack of global pattern aggregation capabilities of the mechanisms that are used. This is especially the fact when dealing with graphs, where nodes struggle to keep track of long range or graph-level dependencies, and tend to be prone to oversmoothing and oversquashing. The domains of generative graph modeling and multi-link prediction can also particularly suffer from this phenomena, as there is no general mechanism that dictates the global distribution of the generated nodes or links, which makes the inference of valid graphs quite challenging. Our goal is to find, for any input set of nodes $V$, a plausible corresponding set of edges $E$. This is often treated as a probabilistic problem that implies estimating the conditional joint distribution of the edges $E$ knowing $V$, which is a classical formulation for conditional generative models. Several variational methods such as [5] or [6] have been proposed. The first one relies on a node-level sampling procedure that assumes nodes’ independence, the second one uses a graph level latent space that is then decoded into a graph. The main difficulty of such a method is that it requires complex matching algorithms to compare the reconstructed nodes and edges with the labeled ones. When the decoding is node-conditioned, the conditioning signal also serves as a node identifier (a reconstructed node to which is concatenated an attribute $a_{k}$ will be compared to the labeled node with attribute $a_{k}$ if the attributes are unique), which can allow avoiding the matching process. Other methods use generative adversarial formulations to generate realistic graphs. In [7], the authors use a unconditional Generative Adversarial Network (GAN) helped by objective-driven Reinforcement Learning, an iterative GAN is proposed in [8] while a temporal sequence GAN is showcased in [9]. Flow-based models, such as autoregressive [10] are also a possible solution for graph generation. More recently, Langevin-esque variational methods have become really popular to generate graphs, in particular Denoising Diffusion Probabilistic Models (DDPM) such as [11] or [12] that respectively provide rotation invariance and discrete, sparsity-encouraging noise schedule. Both use Graph Transformers [13] to denoise at both node and edge levels, as edge attributes are modulated by the nodes’ attention scores, which they also modulate, in an interdependent way. Graph Transformers are quite popular and useful to operate at both node and edge levels, and the attention mechanism is particularly well fitted for graph problems, and grants permutation invariance. As mentioned beforehand, as expressive as it is, the attention mechanism can still struggle to learn graph-level features and is not totally immune to some artifacts, whether by saturating some central nodes with information or creating unwanted high predicted-link-density clusters in areas with an important concentration of nodes. One popular way of improving and providing a form of conditioning to the feature aggregator is to use Feature-wise Linear Modulation (FiLM) [14] [15], which maps a conditional signal to an affine modulation of a feature. For instance, in [12], the authors use FiLM layers to condition the layers on the time step of the diffusion process, to feature some additional graph-level attributes, and encapsulate some global, statistical information about the nodes and edges. In [16], the authors propose using registers, inspired by [17], which consists in adding dummy nodes that can be attended to by the true nodes as supplementary information storage. Using attention on the true tokens to gather general information can also so be seen in other domains such as natural language processing with BERT’s CLS tokens [18], in CrossVIT [19] or in [20] for session-based recommendation. While such method is particularly light, it still has limited expressiveness for graphs and does not provide a flexible mechanism to tune the prediction of the links. Our work will then mainly consist in leveraging flexible and expressive information gathering with cross-attention-based register tokens, and in using this information as a conditioning signal for node-level and edge-level modulations.
In this work we:
  • present our CAM tokens’ architecture in detail
  • provide two different architectures that can typically benefit from CAM tokens
  • present benchmarks that display the merits of our method.
@@ 2 Our problem
Throughout this work, we will tackle a simple toy problem that typically requires both local and global arbitration: given a set of nodes $V$ described by random 2D coordinates, we wish to predict a minimal set of edges $E$, subject to several constraints. One node can not have more than $k$ edges, edges must respect a range constraint, which implies that their length has to be inferior to $d$. The predicted edges have to minimize the number of connected components, hence, if it is possible regarding the range constraint, the graph has to be connected, if not, the number of connected components has to be as small as possible. This can be viewed as a constrained Minimum Spanning Tree problem with all edges’ weights being equal to 1. It then implies respecting some local constraints and requires some higher level perception of connectivity of the graph, which is typically what we try to improve, and what node-level message passing schemes can struggle to tackle. The aim of such an approach is not necessarily to get neural networks to perfectly mimic exact optimization results, but to be able to infer plausible graphs that resemble a dataset of graphs that follow the aforementioned constraints. In order to be valid, the inferred edges $E$ have to: - be individually valid, present similar properties as the dataset edges and, most importantly, respect the range constraints
@@ 5 State of the art generative method: Probabilistic Denoising
Recent works have proven Denoising Diffusion Probabilistic Models (DDPMs) to be particularly effective to tackle the task of generating plausible graphs. In order to approximate the distribution that we assume our data to follow, graph DDPMs propose, in the continuous setting, to search for a $p_{\theta}^{t}(\mathbf{G^{t-1}}|\mathbf{G^{t}})$ that approximates a markovian reverse noise process. This, contrarily to directly maximizing the likelihood of $p_{\theta}(\mathbf{E}|V)$, is directly tractable. In our case, we want to compute the edges $E$ given the nodes $V$, hence we apply the denoising process on edges only. We would then wish to find $p_{\theta}^{t}(\mathbf{E^{t-1}}|\mathbf{E^{t}},V)=p_{\theta}^{t}(\mathbf{E^{t-1}}|\mathbf{G^{t}})$.
We rely on discrete denoising with data-aware state transition matrices, as it proved to be the most effective method in [12]. We then directly model $p_{\theta}^{t}(\mathbf{E}|\mathbf{G^{t}})$. The noise is applied to uniform state transition matrices between the node categorical attributes. Here we consider a binary 1-dimensional problem, this is why our "transition matrix" is only a scalar $m$ that is correlated with the true distribution of the dataset links.
The noise at step $t$ is given by $\bar{Q}^{t}=\bar{\alpha}^{t}I+\bar{\beta}^{t}\mathbf{m}$, where $\bar{\alpha}^{t}=\cos({0.5\pi(t/T+s)}/(1+s))^{2}$ with a small $s$ and $\bar{\beta}^{t}=1-\bar{\alpha}^{t}$.
The learning procedure solely resides in minimizing $\sum_{ij}l_{BCE}(E,p_{\theta,ij}(G^{t}))$.
Sampling with the DDPM to generate new graphs iteratively follows
$E^{t-1}\ \mathtt{\sim}\ \prod_{ij}\hat{p}^{E}_{(\theta,t,ij)}(G^{t})\>q(e_{ij}^{t-1}|e_{ij}=e,e_{ij}^{t}),$ (17)
with $\hat{p_{\theta}}$ being an estimation of $p$ by a neural network parametrized by $\theta$, $q$ representing the distribution of the scheduled noisy transition matrix.